OpenStudy (anonymous):

determine wether each rule represents and exponential function. y=12*x^2 y=7x+3

OpenStudy (anonymous):

@AdamK

OpenStudy (solomonzelman):

exponential function is when the variable in the exponent.

OpenStudy (solomonzelman):

Such as \(\large\color{slate}{ y=e^x }\)

OpenStudy (solomonzelman):

this is exponential or not, what do you think ?

OpenStudy (anonymous):

the first on is the second is not

OpenStudy (solomonzelman):

the second one is not, correct. The first one is, are you sure?

OpenStudy (solomonzelman):

\(\large\color{slate}{ y=x^2 }\) is just a parabola, and \(\large\color{slate}{ y=12x^2 }\) is a parabola increased by the scale factor of 12.

OpenStudy (solomonzelman):

But, exponential function is something like \(\large\color{slate}{ y=a(b)^x}\) where \(\large\color{black}{ \huge{ \begin{array}{| l | c | r |} \hline \scr~~~x~~ & \scr y \\ \hline \scr~~~0~ & \scr a \\ \hline \scr~~~1~ & \scr a\cdot b \\ \hline \scr~~~2~ & \scr a\cdot b\cdot b \\ \hline \scr~~~3~ & \scr a\cdot b\cdot b \cdot b \\ \hline \end{array} } }\) and on.... multiplying times b every time x goes up by 1. (where b>1)

OpenStudy (solomonzelman):

I mean where b>0

OpenStudy (solomonzelman):

but in case of y=x^2 you are just taking perfect squares x y 1 1 2 4 3 9 4 16 when you have y=12x^2, then x y 1 12 2 48 3 108 4 196 and on... where all you are doing is mltiplying all the perfect squares times 12.

OpenStudy (triciaal):

the first one is an exponential function the variable is raised to a power the change is "y" is not linear